OpenStudy (anonymous):

simplify 2^3 x^3 y^-2*2^-5 xy^2

OpenStudy (anonymous):

(3x^2y)(5xy^3) = (3)(x^2)(y)(5)(x)(y^3) = (3)(5)(x^2)(x)(y)(y^3) = 15x^3y^4